Two flasks are connected with a stopcock. Flask #1 has a volume of 2.5 L and contains oxygen gas at a pressure of 0.70 atm. Flask #2 has a volume of 3.8 L and contains hydrogen gas at a pressure of 1.25 atm. When the stopcock between the two flasks is opened and the gases are allowed to mix, what will the resulting pressure of the gas mixture be?



1
Expert's answer
2021-05-21T02:33:06-0400

"P_1V_1+P_2V_2 = P_{mix}V_{mix}"


"0.75 \times 5+1.25 \times 8 = P_{mix}(5+8)"

"P_{mix} = 1.057"